# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v3

@@ V-B4 Adversarial Perturbation-Based defenses
WF Attack Model
Attack Rate
Acc (%)
Defense Rate
[…]
1) Protection against multiple-round intersection attacks; 2) Randomized target perturbation strategy; 3) Less information leakage.
1) Deployment complexity; 2) Requires prior knowledge of the complete traffic burst patterns.
@@
1) High evasion rate and low overhead efficiency; 2) Only a few past bursts of a trace are needed.
1) Vulnerability to adversarial training; 2) Time consuming in searching fuzzy adversarial perturbation direction; 3) Only modifies the future part of the traffic trace, limiting the defense’s effectiveness.
@@
1) Works in real-time on live traffic; 2) High transferability and flexibility.
1) Vulnerability to defense-aware retraining; 2) Potential to informative features leakage.
[…]
1) Prior knowledge of traffic patterns is not required; 2) Defense against adversarial training.
1) Fixed-length UAP matching may reduce precision in real-world traces; 2) Improperly tuning may increase bandwidth overhead with loss-tuning-based bandwidth control.
@@
Minipatch [li2022minipatch] AWF [rimmer2017automated] NA >>99.70
1) Real-time and real-world deployment; 2) Lower bandwidth overhead with few dummy packets injection.
1) Unable to withstand attacks that involve retraining the classifier; 2) Requires prior knowledge of visited website.
@@
1) Robust against adversarial training; 2) Universal and fast perturbation generation; 3) Efficient and lightweight.
1) Trade-off tuning required; 2) No real-time dynamic adaptation; 3) Potential for overhead increase with more data.
@@
WF Attack Model
Attack Rate
Acc (%)
Defense Rate
[…]
1) No prior knowledge required; 2) Defend against adversarial training by random response; 3) Can mimic realistic traffic patterns of different webpages by GAN.
1) High bandwidth and latency overhead, and not cost-effective for real-world use; 2) Feature leakage still possible.
@@
1) Low communication overhead; 2) GAN-based universal perturbation generator and pluggable transport compatible.
1) Defense efficiency depends on substitute classifier; 2) Potential to informative features leakage.
@@
Low bandwidth overhead and real-time implementation.
1) Requires prior website knowledge; 2) Performance depends on surrogate model quality; 3) Sensitive to network condition.
@@
Gong et al. [gong2022surakav] proposed a solution using Generative Adversarial Networks (GANs) to generate realistic traffic patterns that mimic the behavior of various web pages. These synthetic traces are used to regulate packet transmission, obscuring the website being visited by introducing randomness in both packet sizes and sending times. This dynamic adjustment of packet bursts based on buffered data makes it significantly harder for attackers to distinguish between different websites. Jiang et al. [jiang2022effective] introduced Targeted LightwEight Defense (TED), a lightweight and efficient method to defend against WF attacks in the Tor network. TED introduces trace-level defenses, customizing the defense for each unique trace—even for the same website. In addition, TED leverages a fully convolutional network (FCN) combined with adapted Grad-CAM to extract key discriminative features from traffic traces. These are the elements deep learning-based WF attacks rely on. TED then precisely perturbs these key features by injecting a small number of dummy packets, misleading the attacker’s classifier without heavy overhead.
@@ V-B5 Other WF defense methods
Gu et al. [gu2023online] presented an innovative and practical solution for defending against WF attacks, showcasing the potential of adversarial techniques in enhancing online privacy without incurring prohibitive costs or overhead. This research contributes significantly to the field of network security, specifically in protecting user privacy while browsing the web. Liu et al. [liu2023smart] presented SMART, a lightweight defense that splits traffic across multiple Tor entry relays, disrupting statistical features used by WFAs. SMART offers a practical, efficient defense against WFAs by combining multi-path routing and redundancy. However, Tor enables anonymous communication but remains vulnerable to flow correlation attacks that match traffic patterns between its entry and exit points. MUFFLER [seo2025muffler] addresses these issues by dynamically reshaping egress traffic through real-time connection mapping without adding padding or delays. It effectively disrupts correlation attacks, achieving only 1% true positive rate at 1% false positive rate, with minimal bandwidth and lower latency—while maintaining compatibility with the existing Tor infrastructure.
@@ VI-A1 Webpage assumption
Static webpage. The static webpage assumption suggests avoiding WF attacks on sites where content changes frequently, such as news websites or video content platforms that update every few minutes [wang2013improved]. Training a traffic classifier is often time-consuming, and when webpage content changes, the attacker must retrain the classifier using new browsing traffic, which is very costly. As a result, attacks typically target static webpages or websites with regularly updated content, such as government agency sites.
@@
VI-A2 Concept drift
Concept drift in WF attacks refers to the phenomenon where changes in webpage content over time alter the associated traffic patterns, thereby reducing the accuracy of classification models. Modern websites often update their content frequently, whether through dynamic elements, advertisements, multimedia, or layout modifications, all of which can significantly shift traffic characteristics. Training a traffic classifier is typically resource-intensive, requiring large volumes of labeled data and extensive computation.
@@ VI-A2 Concept drift
When concept drift occurs, attackers are forced to retrain classifiers on newly collected traces—a process that is both costly and time-consuming. Many WF attacks focus on relatively static websites or those with predictable, infrequent updates (e.g., government portals) to maintain effectiveness. However, this assumption of stable targets highlights a critical limitation of WF attacks in dynamic, real-world environments. It underscores the need for adaptive models that can handle evolving traffic patterns without requiring complete retraining.
@@ VI-A4 Network situation
In WF attack and defense studies, researchers typically assume that the collected datasets are captured under consistent network conditions. However, network congestion can introduce significant delays, which directly affect traffic patterns. WF models often rely on traffic features extracted within fixed time slots, as the exact start and end of a website trace are usually unknown. Prolonged delays increase the interval between consecutive packets and reduce the number of packets captured within the designated time slot. Consequently, the extracted features may no longer accurately represent the website’s traffic pattern, leading to degraded attack performance. This effect is particularly pronounced in dynamic or congested network environments, highlighting the need for WF models and defenses to account for variable latency and packet timing variability.
@@ VI-B2 WF attack model building
Beyond these surface-level features, future research can explore the interaction processes between communication components—such as client-server exchanges—and enrich traffic data with more semantic information from the application layer. Such an approach can offer deeper insights into the fundamental reasons behind traffic differences across classes. Additionally, suitable evaluation metrics are required to quantitatively assess the effectiveness of various representations.
@@
Robustness. Traffic representations should be resilient to changes in Tor traffic without requiring substantial modifications to their initial design. Such changes may take various forms, including protocol upgrades, content updates, and fluctuations in network conditions. Statistical representations tend to be highly sensitive to these changes, as even minor variations can render the original representations ineffective. This sensitivity is evident in the quantitative measures of information leakage for different statistical features—features that are most informative for website fingerprinting often vary drastically when the traffic is perturbed.
@@ VI-C2 Unreliable Evaluations of Defense Schemes
VI-C3 Countering Traffic Analysis Attacks in Low-Latency Anonymity Networks
@@
  • Theoretical Perspective: To build more robust defenses, it is essential to strengthen theoretical foundations—particularly those grounded in deep learning itself. One promising direction is to demystify the black-box nature of deep learning, which can facilitate the development of more fine-grained and targeted defense strategies.
  • Practical Perspective: Defenses against traffic analysis attacks in anonymity networks can be implemented at multiple levels, namely the network, protocol, and application layers—offering flexibility and the potential for layered protection.
@@ VI-D Future works
There is still a gap between current research and real-world deployment. As discussed before, the field faces several limitations, including overly strong assumptions, severe concept drift issues, and insufficient data. Addressing these issues is crucial for improving the practicality and reliability of website fingerprinting models in real-world environments.
@@ VI-D1 Relaxing Identification Assumptions
First, relaxing identification assumptions is a key direction for future website fingerprinting attack research. By reducing assumptions about user behavior, attacker capabilities, and webpage characteristics, while taking more real-world factors into account, the practicality and robustness of website fingerprinting models can be improved to better meet real-world application needs. Second, for user behavior assumptions, future research could focus on designing more flexible user behavior models. These models should allow for realistic browsing behaviors, such as opening multiple browser tabs simultaneously and performing background activities, to more accurately simulate real user behavior. Third, for attacker behavior assumptions, future studies can explore more diverse and complex fingerprinting attack strategies, potentially incorporating simulation-based training in realistic user environments. Fourth, for webpage assumptions, future work could expand to include a wider variety of page types—such as internal pages, frequently updated pages, noisy pages, and cached pages—thereby enhancing the model’s adaptability and analytical capabilities across different types of web content.
VI-D2 Mitigating the Concept Drift Problem
In WF models, classifiers are typically trained on pre-collected datasets. However, as time passes, changes in website traffic patterns are not reflected in the trained classifier, leading to a significant drop in classification accuracy. Most existing studies have focused on using deep learning to extract more general and stable fingerprinting features to counteract concept drift, but these methods have reached a performance bottleneck.
Future research can explore model and data update techniques to enable continuous improvement of classifiers. For model updates, approaches such as incremental learning or online learning can be used to incorporate the latest website traffic traces into the training process without retraining the entire model. This improves efficiency and meets real-time update requirements. For data updates, fingerprint data can be continuously monitored and collected. Techniques such as sliding windows can be used to partition the data and regularly discard outdated batches to keep the dataset current. Furthermore, ensemble learning methods can be introduced to build multiple diverse models, reducing the bias and variance of individual neural networks. This can enhance model robustness and improve performance on data affected by concept drift.
@@ VI-D6 Large Language Models (LLMs) in WF attacks and defenses
LLMs are advanced deep learning models trained on massive text corpora to understand and generate natural language. Built on transformer architectures, LLMs capture complex contextual and sequential dependencies, enabling them to generalize across diverse data modalities beyond text. Recently, researchers have explored applying LLMs to WF [jiao2025llm, song2025redefining], where encrypted traffic traces are represented as symbolic sequences. In the LLM-based paradigm, encrypted traffic traces are converted into symbolic or textual sequences that capture packet-level or burst-level characteristics. LLMs can be integrated into WF tasks in multiple ways:

### v4

@@ V-B4 Adversarial Perturbation-Based defenses
WF Attack Model Attack Rate Acc (%) Defense Rate (%)
@@
[…]
Minipatch [li2022minipatch] AWF [rimmer2017automated] NA >>99.70 1) Real-time and real-world deployment; 2) Lower bandwidth overhead with few dummy packets injection. 1) Unable to withstand attacks that involve retraining the classifier; 2) Requires prior knowledge of visited website.
@@
WF Attack Model Attack Rate Acc (%) Defense Rate (%)
@@
[…]
@@
Gong et al. [gong2022surakav] proposed a solution using Generative Adversarial Networks (GANs) to generate realistic traffic patterns that mimic the behavior of various web pages. These synthetic traces are used to regulate packet transmission, obscuring the website being visited by introducing randomness in both packet sizes and sending times. This dynamic adjustment of packet bursts based on buffered data makes it significantly harder for attackers to distinguish between different websites. Jiang et al. [jiang2022effective] introduced Targeted LightwEight Defense (TED), a lightweight and efficient method to defend against WF attacks in the Tor network. TED introduces trace-level defenses, customizing the defense for each unique trace, even for the same website. In addition, TED leverages a fully convolutional network (FCN) combined with adapted Grad-CAM to extract key discriminative features from traffic traces. These are the elements deep learning-based WF attacks rely on. TED then precisely perturbs these key features by injecting a small number of dummy packets, misleading the attacker’s classifier without heavy overhead.
@@
In 2025, Xie et al. [xie2025gapdis] introduced GAPDiS, a gradient-based perturbation for WF defense. To address the Gradient Incompatibility Problem, the authors introduce an offset vector and define its cosine similarity with the gradient as a reward signal. To efficiently generate universal perturbations that are effective across all trained website categories, GAPDiS integrates the global exploration capability of tabu search with gradient-assisted local optimization. Experimental results on the AWF dataset demonstrate that GAPDiS reduces WF classification accuracy from over 98% to below 7% with only 2.56% bandwidth overhead, yielding a 68.1 improvement over SOTA defenses.
@@ V-B5 Other WF defense methods
Gu et al. [gu2023online] presented an innovative and practical solution for defending against WF attacks, showcasing the potential of adversarial techniques in enhancing online privacy without incurring prohibitive costs or overhead. This research contributes significantly to the field of network security, specifically in protecting user privacy while browsing the web. Liu et al. [liu2023smart] presented SMART, a lightweight defense that splits traffic across multiple Tor entry relays, disrupting statistical features used by WFAs. SMART offers a practical, efficient defense against WFAs by combining multi-path routing and redundancy. However, Tor enables anonymous communication but remains vulnerable to flow correlation attacks that match traffic patterns between its entry and exit points. MUFFLER [seo2025muffler] addresses these issues by dynamically reshaping egress traffic through real-time connection mapping without adding padding or delays. It effectively disrupts correlation attacks, achieving only 1% true positive rate at 1% false positive rate, with minimal bandwidth and lower latency-while maintaining compatibility with the existing Tor infrastructure.
@@ VI-A1 Webpage Assumption
Static webpage. The static webpage assumption suggests avoiding WF attacks on sites where content changes frequently, such as news websites or video content platforms that update every few minutes [wang2013improved]. Training a traffic classifier is often time-consuming, and when webpage content changes, the attacker must retrain the classifier using new browsing traffic, which is very costly. As a result, attacks typically target static webpages or websites with regularly updated content.
@@
VI-A2 Model Drift
Model drift refers to the gradual degradation of an ML/DL model’s performance over time. This degradation can arise from changes in the data distribution, shifts in the model’s objectives, or variations in the operating environment. Model drift is commonly categorized into two main types: data drift and concept drift. We observe that many prior studies conflate these two concepts; therefore, to avoid such confusion in future work, we provide precise definitions of each.
Data drift. In ML/DL, data drift (also known as covariate shift) occurs when the statistical distribution of input data changes over time, creating a mismatch between the environment in which a model was trained and the live environment where it is deployed. Because models are optimized based on the patterns of the training set, this divergence often leads to “model decay”, where predictive performance and reliability degrade significantly [moreno2012unifying]. Importantly, under data drift, the underlying relationship between inputs and labels remains unchanged.
A mathematical representation of data drift can be expressed as follows:
$p_{train}(X)\neq p_{test}(X),\ p(y\mid X)\ \text{unchanged}$ (12)
Where $p(X)$ denotes the probability distribution of input features, and $p(y\mid X)$ is the conditional probability of the target given the features (the model’s “logic”).
Data drift in WF attacks refers to the phenomenon where changes in webpage content over time alter the associated traffic patterns, thereby reducing the accuracy of classification models. Modern websites often update their content frequently, whether through dynamic elements, advertisements, multimedia, or layout modifications, all of which can significantly shift traffic characteristics. Despite these changes, the semantic label of the traffic, namely, the visited website, remains unchanged.
We observe that many existing WF datasets use the domain name as the classification label. Even when subdomains or sub-URLs are added, removed, or modified, they are typically mapped to the same domain-level label. In such cases, although the input traffic distribution shifts, the label definition does not change. Therefore, this phenomenon should be properly characterized as data drift rather than concept drift. Misclassifying this scenario as concept drift can lead to inappropriate modeling assumptions and evaluation strategies. Accurately distinguishing data drift from concept drift is thus essential for designing robust WF defenses and attacks, as well as for ensuring fair and meaningful experimental evaluations.
Concept drift. In contrast to data drift, concept drift specifically refers to a shift in the underlying relationship between input features and the target variable. In these scenarios, the statistical properties of the target variable itself change, meaning that even if the input data remains consistent, the model’s predictions become increasingly inaccurate because the “ground truth” logic it once learned is no longer valid [gama2014survey].
Formally, concept drift between time points $t_{0}$ and time point $t_{1}$ can be defined as:
$\exists X:p_{t_{0}}(X,y)\neq p_{t_{1}}(X,y)$ (13)
where $p_{t_{0}}(X,y)$ denotes the joint distribution of the input variables $X$ and the target variable $y$ at time $t_{0}$.Such changes can be characterized by variations in different components of this distribution: (1) shifts in the class prior probabilities $p(y)$; 2) changes in the class-conditional distributions $p(X\mid y)$; and consequently, 3) changes in the posterior probabilities $p(y\mid X)$, which directly affect model predictions.
Concept drift in WF attacks refers to the phenomenon where changes in website semantics or functionality alter the relationship between network traffic features and the corresponding website labels. Unlike data drift, where traffic patterns change while the label definition remains the same, concept drift arises when website redesigns, structural overhauls, protocol changes, or fundamental modifications to page logic significantly reshape how content is delivered and interacted with. As a result, previously learned mappings between traffic characteristics and website identities become invalid, even if the domain label remains unchanged. Moreover, modern websites increasingly incorporate content, scripts, and multimedia from external domains, further modifying traffic generation behavior. These factors collectively invalidate previously learned mappings between traffic patterns and website identities, even when the domain label itself remains unchanged. Consequently, concept drift leads to significant performance degradation and necessitates model retraining or continual adaptation to sustain effective WF attacks.
@@ VI-A2 Model Drift
When concept drift occurs, attackers are forced to retrain classifiers on newly collected traces, a process that is both costly and time-consuming. Many WF attacks focus on relatively static websites or those with predictable, infrequent updates to maintain effectiveness. However, this assumption of stable targets highlights a critical limitation of WF attacks in dynamic, real-world environments. It underscores the need for adaptive models that can handle evolving traffic patterns without requiring complete retraining.
@@ VI-A4 Network Situation
In WF attack and defense studies, researchers typically assume that the collected datasets are captured under consistent network conditions. However, network congestion can introduce significant delays, which directly affect traffic patterns [shen2025swallow]. WF models often rely on traffic features extracted within fixed time slots, as the exact start and end of a website trace are usually unknown. Prolonged delays increase the interval between consecutive packets and reduce the number of packets captured within the designated time slot. Consequently, the extracted features may no longer accurately represent the website’s traffic pattern, leading to degraded attack performance. This effect is particularly pronounced in dynamic or congested network environments, highlighting the need for WF models and defenses to account for variable latency and packet timing variability.
@@ VI-B2 WF Attack Model Building
Beyond these surface-level features, future research can explore the interaction processes between communication components, such as client-server exchanges, and enrich traffic data with more semantic information from the application layer. Such an approach can offer deeper insights into the fundamental reasons behind traffic differences across classes. Additionally, suitable evaluation metrics are required to quantitatively assess the effectiveness of various representations.
@@
Robustness. Traffic representations should be resilient to changes in Tor traffic without requiring substantial modifications to their initial design. Such changes may take various forms, including protocol upgrades, content updates, and fluctuations in network conditions. Statistical representations tend to be highly sensitive to these changes, as even minor variations can render the original representations ineffective. This sensitivity is evident in the quantitative measures of information leakage for different statistical features-features that are most informative for website fingerprinting often vary drastically when the traffic is perturbed.
@@ VI-C2 Unreliable Evaluations of Defense Schemes
VI-C3 Deployment in Real-World Environments
Existing WF defense mechanisms often rely on computationally intensive models, requiring high-performance hardware or specialized accelerators to achieve acceptable performance. Such requirements are difficult to satisfy in real-world settings, particularly on resource-constrained devices, especially for end-user devices. Moreover, the overhead introduced by these defenses can increase latency and energy consumption, negatively impacting user experience. As a result, despite their effectiveness in controlled environments, many proposed WF defenses face significant challenges in terms of scalability, practicality, and widespread deployment. Addressing these limitations remains critical for translating WF defense research into deployable, real-world solutions.
VI-C4 Countering Traffic Analysis Attacks in Low-Latency Anonymity Networks
@@
  • Theoretical Perspective: To build more robust defenses, it is essential to strengthen theoretical foundations-particularly those grounded in deep learning itself. One promising direction is to demystify the black-box nature of deep learning, which can facilitate the development of more fine-grained and targeted defense strategies.
  • Practical Perspective: Defenses against traffic analysis attacks in anonymity networks can be implemented at multiple levels, namely the network, protocol, and application layers-offering flexibility and the potential for layered protection.
VI-C5 Lack of Adversarial Training Evaluation
When a WF defense mechanism is deployed, any traffic collected by an adversary would reflect the defended Tor traffic. Consequently, the adversary’s model would be trained on this altered traffic. Under such conditions, it is crucial to evaluate the model’s performance under adversarial training, as it provides a more realistic measure of the defense’s effectiveness. However, most existing studies focus solely on the defense’s performance metrics and neglect to assess the impact on models trained adversarially, potentially overestimating the practical effectiveness of the proposed defenses.
@@ VI-D Future works
A gap still exists between current research and real-world deployment. As discussed before, the field faces several limitations, including overly strong assumptions, severe concept drift issues, and insufficient data. Addressing these issues is crucial for improving the practicality and reliability of website fingerprinting models in real-world environments.
@@ VI-D1 Relaxing Identification Assumptions
First, relaxing identification assumptions is a key direction for future WF attack research. By reducing assumptions about user behavior, attacker capabilities, and webpage characteristics, while taking more real-world factors into account, the practicality and robustness of WF models can be enhanced to better meet the needs of real-world applications. Second, for user behavior assumptions, future research could focus on designing more flexible user behavior models. These models should allow for realistic browsing behaviors, such as opening multiple browser tabs simultaneously and performing background activities, to more accurately simulate real user behavior. Third, for attacker behavior assumptions, future studies can explore more diverse and complex fingerprinting attack strategies, potentially incorporating simulation-based training in realistic user environments. Fourth, for webpage assumptions, future work could expand to include a wider variety of page types, such as internal pages, frequently updated pages, noisy pages, and cached pages, thereby enhancing the model’s adaptability and analytical capabilities across different types of web content.
VI-D2 Mitigating the Model Drift Problem
In WF models, classifiers are typically trained on pre-collected datasets. However, as time passes, changes in website traffic patterns are not reflected in the trained classifier, leading to a significant drop in classification accuracy. Most existing studies have focused on using DL to extract more general and stable fingerprinting features to counteract model drift, but these methods have reached a performance bottleneck.
Future research can explore model and data update techniques to enable continuous improvement of classifiers. For model updates, approaches such as incremental learning or online learning can be used to incorporate the latest website traffic traces into the training process without retraining the entire model. This improves efficiency and meets real-time update requirements. For data updates, fingerprint data can be continuously monitored and collected. Techniques such as sliding windows can be used to partition the data and regularly discard outdated batches to keep the dataset current. Furthermore, ensemble learning methods can be introduced to build multiple diverse models, reducing the bias and variance of individual neural networks. This can enhance model robustness and improve performance on data affected by model drift.
@@ VI-D6 Large Language Models (LLMs) in WF Attacks and Defenses
LLMs are advanced DL models trained on massive text corpora to understand and generate natural language. Built on transformer architectures, LLMs capture complex contextual and sequential dependencies, enabling them to generalize across diverse data modalities beyond text. Recently, researchers have explored applying LLMs to WF [jiao2025llm, song2025redefining], where encrypted traffic traces are represented as symbolic sequences [zhao2025embracing, long2025survey]. In the LLM-based paradigm, encrypted traffic traces are converted into symbolic or textual sequences that capture packet-level or burst-level characteristics. LLMs can be integrated into WF tasks in multiple ways:
@@
Recent advances in time-series language models (TSLMs) [jin2023time, wang2024timemixer, shi2024time, langer2025opentslm] have the potential to substantially impact both WF attacks and defenses. In a WF attack, an adversary observes packet-level metadata, such as packet direction, size, and bursts, and represents this information as a time series. TSLMs, particularly Transformer-based architectures, process these sequences analogously to natural language: packets or traffic bursts act as tokens, and an entire page load corresponds to a document. Through self-attention, these models can capture long-range dependencies across the full traffic trace, enabling them to learn high-level structural patterns of website loading behavior that are difficult to conceal through encryption alone.
Defending against WF attacks empowered by time-series language models is especially challenging, as many traditional defenses rely on static or low-level traffic transformations, such as injecting dummy packets or shaping transmission rates. Although these techniques can obfuscate traffic features used by simpler classifiers, TSLMs can often learn to discount such regularities and instead exploit residual structural information. As a result, recent defenses have shifted toward adaptive and learning-aware strategies designed to disrupt high-level temporal correlations. Examples include randomizing page load phases, mixing traffic from multiple sources, and applying adversarial traffic shaping that explicitly accounts for the attacker’s model.